What is the hydroxide-ion concentration of an aqueous solution having a pH of 4.72?
1
Expert's answer
2020-05-15T10:05:23-0400

Solution:

pH = 4.72

The hydronium ion concentration can be found from the pH:

pH = − log [H3O+]

Thus:

[H3O+] = 10−pH

[H3O+] = 10−4.72 = 1.906×10−5

[H3O+] = 1.906×10−5 M


The hydroxide ion concentration can be calculated using the Kw constant:

Kw = [H3O+]×[OH] = 1.0×10-14

Therefore,

[OH] = Kw / [H3O+]

[OH] = (1.0×10−14) / (1.906×10−5) = 5.246×10−10

[OH] = 5.25×10−10 M


Answer: [OH] = 5.25×10−10 M